OpenStudy (anonymous):

plz solve this one The K.E of a satellite in an orbit close to the surface of earth is K.it's maximum KE so as to escape from the gravitational field of earth is _________ .

OpenStudy (irishboy123):

the gravitational potential energy U1 of the satellite near earth will be U1 = -GMm/r (obtained from Newton's inverse square la for gravitation), with r = radius earth. thus total energy E1 near earth = K + U1. if E2 = energy at r = infinity (well, far enough away for PE --> 0), then E2 = 0 as U2 = 0 (0 because no gravitational potential as you have moved away from earth), K2 = 0 (0 again 0 because you only had enough energy to escape the earths field in the first place). thus E2 = E1 = K + U1 = 0 thus K = GMm/r

OpenStudy (anonymous):

answer is 2K

OpenStudy (irishboy123):

In orbit: mv^2/ r = GMm/ r^2 (balances forces, right?) It follows that: K1 (kinetic energy satellite has in orbit ) = 1/2 mv^2 = (1/2) * GMm/r We conclude that K (energy it needs to escape) = 2 K1 Note the factor of 1/2 in previous post. I can write that out more sensibly if it matters.

OpenStudy (anonymous):

no need to to that now :)